You are here

Beam delivery update: Short maintenance and beam development tomorrow

header

08 December 2014

The cyclotron is currently delivering ~110 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. This will continue until ~07:00 tomorrow, at which point the cyclotron will go off for a short maintenance period followed by beam development. This will mark the end of BL1A operation for 2014. Normal operation will resume ~20:30 tomorrow with BL1B, BL2A, and BL2C4 (as PIF is expected to run only on BL1B for the week).

At ISAC, delivery to GRIFFIN ended this morning. Beam was tuned to the ISAC implantation station for experiment L112 earlier today and the mass separator room accessed to install a target foil for that experiment; 213Fr delivery is being set up as I type this. Assuming tonight’s implantation is successful, we’ll begin setting up for FrPNC experiment S1218 tomorrow evening for delivery on Wednesday. TRILIS was also set up for Mg ionization this afternoon; the yields of 27–32Mg were measured and found to be two or more orders of magnitude less than what we’d normally expect from a UCx target. This is in line with what we've observed for other elements with the target in its current, compromised configuration. Based on these yields, I expect that the RIB schedule will remain as it currently stands for the rest of the year.

Also at ISAC, 76Se delivery from the OLIS multicharge ion source to DRAGON for measurements of Se charge-state distributions will continue tomorrow. That should wrap up scheduled OLIS delivery and ISAC accelerator operation for the year.

-- ACM